OpenStudy (chaise):

(a+bi)^2 = -2+5i; a,b>o

OpenStudy (anonymous):

Again, foil it out. If you see an \(i^2\) swap it with a (-1)

OpenStudy (chaise):

\[(a+bi)^{2} =-2+5i\]

OpenStudy (anonymous):

\[(a + bi)(a+bi) \]\[= a^2 + 2abi + b^2i^2 \]\[=(a^2 - b^2) + (2ab)i = -2+5i\] \[\implies \begin{array}{ccc} a^2 - b^2 = -2 & and & 2ab = 5\end{array}\]

OpenStudy (chaise):

Does this now become some sort of simultaneous equation?

OpenStudy (anonymous):

It is a system of equations that you'll have to solve for a and b.

OpenStudy (anonymous):

I suggest the substitution method..

OpenStudy (anonymous):

\[b = \frac{5}{2a} \implies b^2 = \frac{25}{4a^2}\]

OpenStudy (chaise):

You're a legend. Thanks a lot! A lot easier then I was told in class.

OpenStudy (anonymous):

How was it explained in class?

OpenStudy (anonymous):

The only real trick is to foil, then equate the coefficients of the real and imaginary parts.

OpenStudy (anonymous):

Is this algebra 2?

OpenStudy (chaise):

It was explained VERY differently. Using the quadratic formula, amung other things. I don't know - we call it Maths C in Australia. Complex numbers, as you may have guessed.

OpenStudy (anonymous):

Well you will need the quadratic formula

OpenStudy (anonymous):

Once you substitute in the \(b^2\) term

OpenStudy (chaise):

Cool. Thanks sir.

OpenStudy (anonymous):

Oh, an aussi eh? We don't get many of them around here.

OpenStudy (anonymous):

This is for pre-university though yes?

OpenStudy (chaise):

I guess so. I'm in year 11. I go to university after next year. Why do you ask?

OpenStudy (anonymous):

Just curious

OpenStudy (chaise):

May I ask your profession?

OpenStudy (anonymous):

I'm a student at university. I also do some programming and sysadmin work on the side.

OpenStudy (chaise):

That's good. :)

OpenStudy (anonymous):

So did you get this system solved yet?

OpenStudy (anonymous):

Or are you still working on it?

OpenStudy (chaise):

a^2 -(25/4a^2) = -2 Multiply every term by 4a^2 to get rid of the denominator in the fraction. 4a^4 - 25 = -8a^2 4a^4 + 8a^2 -25 = 0 4x^2 + 8x - 25 = 0 where x = a^2 x = -3.7 x = 1.7 x = a^2 a = sqrt(-3.7) b = sqrt(1.7) a = 1.9i a = -1.9i b = 1.3 b = -1.3 Does this seem about right? Thanks in advance.

OpenStudy (anonymous):

Recall however that the coefficients a and b are both real, and they are both given to be positive.

OpenStudy (anonymous):

So the only valid solution for a is sqrt(1.7)

OpenStudy (anonymous):

And that would make \[b = \frac{5}{2\sqrt{1.7}}\]

OpenStudy (chaise):

Aaaah; because the restrictions on the equation at the start where a,b > 0? So the real answers would be: a = 1.9i b = 1.3 Correct?

OpenStudy (anonymous):

a cannot be 1.9i

OpenStudy (anonymous):

I think you are confusing yourself. back up to the solutions you got for x.

OpenStudy (chaise):

x = a^2 a = sqrt(-3.7) b = sqrt(1.7) a = 1.9i b = 1.3 Square root the negative number you get a result which is i?

OpenStudy (chaise):

How would you go about solving this equation? 4a^4 + 8a^2 -25 = 0

OpenStudy (anonymous):

\[x \in \{-3.7,1.7\}\]\[\implies a^2 \in \{-3.7,1.7\}\]

OpenStudy (anonymous):

You solved it the right way. You just miss-interpreted the results.

OpenStudy (anonymous):

We get two different solutions for a^2

OpenStudy (anonymous):

-3.7 and 1.7

OpenStudy (anonymous):

But a must be real (its the coefficient)

OpenStudy (anonymous):

So the square of a cannot be negative.

OpenStudy (anonymous):

So we can discard the -3.7 result. and we are left with \[a = \pm \sqrt{1.7}\]

OpenStudy (anonymous):

But since we are given that a must be positive we can throw away the negative square root also.

OpenStudy (anonymous):

So for sure \[a = \sqrt{1.7} \approx 1.303\]

OpenStudy (anonymous):

Now we just go back to 2ab = 5 to solve for b.

OpenStudy (anonymous):

The coefficients of complex numbers are always real.

OpenStudy (chaise):

Aaah okay. Thanks so much. We can actually leave it as a^2-b^2 = 2 So if a^2 = 1.7 a^2-b^2 = 1.7 1.7 - b^2 = 2 -b^2 = 0.3 -b = sqrt(0.3) b = -0.54? Or must you use the other equation? the second equation?

OpenStudy (anonymous):

That's not right.

OpenStudy (anonymous):

\[a^2 - b^2 = -2\]\[\implies 1.7 - b^2 = -2 \]\[\implies -b^2 = -3.7\]\[ \implies b^2 = 3.7 \implies b = \sqrt{3.7}\]

OpenStudy (chaise):

Cool. That makes sense! Thank you so much! You time is truely appreciated!

OpenStudy (anonymous):

Sure thing =)

OpenStudy (chaise):

Surely you have something better to do with your time?

OpenStudy (anonymous):

At this moment, it's either this, or sleeping. I'll go with this for a while.